1. a) A fidget spinner is rotating at a certain angular velocity. As it began to slow down, it
took 45 seconds to stop with an average angular acceleration of 0.69 rads-2
. Calculate
the
(i) initial angular velocity in rpm.
(ii) number of revolution the spinner makes before it stops.
1
Expert's answer
2020-11-02T09:23:06-0500

The angular acceleration is


"\alpha=\frac{\omega_2-\omega_1}{t},\\\space\\
\omega_2=\alpha t+\omega_1=0.69\cdot45+0=45.69\text{ rad/s}."

The number of revolutions before it stopped is


"n=\frac{\omega_2}{2\pi}=\frac{45.69}{2\pi}=7.272."